OpenStudy (anonymous):

twice the difference of a number and 11

OpenStudy (anonymous):

Can be clarify your question??

OpenStudy (anonymous):

2(x-11)

OpenStudy (anonymous):

oh !

OpenStudy (anonymous):

thanks girly (:

OpenStudy (anonymous):

No problem. :)

OpenStudy (anonymous):

Sove: 4y-1/10=3y=4/5 please tell me. thanks

OpenStudy (anonymous):

4y - 1/10 = 3y + 4/5 y -1/10 = 4/5 subtract 3y from each side y - 1/10= 8/10 common denominator y = 9/10 add 1/10 to each side

OpenStudy (anonymous):

can you please show me the pic if you have any?

OpenStudy (anonymous):

Never mind it's ok i got it

OpenStudy (anonymous):

You sure?

OpenStudy (anonymous):

name the additive inverse and multiplicative inverse for -4/3

OpenStudy (anonymous):

yesss girlyy,, thanks